Find the sum of vector A and C
A=8i + 8j
C=8i - 16j
1
Expert's answer
2020-02-25T10:44:37-0500

Vectors are added component-wise, i.e. "(x_1 i + y_1 j) + (x_2 i + y_2 j) = (x_1 + x_2) i + (y_1 + y_2) j".

Hence, "(8 i + 8 j) + (8 i - 16 j) = (8 + 8) i + (8 - 16) j = 16 i - 8 j"